# Nonlinear Schrödinger-Poisson systems in dimension two: the zero mass case

### v1

@@ 3. The variational framework
The aim of this section is to show that $\mathcal{I}$ is indeed well-defined and regular in the space $D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$ described in Section 2. First, we state an identity which will play a crucial rôle throughout the paper:
@@
$(u,v)\mapsto A_{0}(u,v):=A_{1}(u,v)-A_{2}(u,v)=\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln|x-y|u(x)v(y)\,{\rm d}x\,{\rm d}y$
Since $b>1$ one has
(3.3) $\ln(b+|x-y|)\leq\ln(b+|x|+|y|)\leq\ln(b+b|x|+b|y|)\leq\ln((b+|x|)(b+|y))=\ln(b+|x|)+\ln(b+|y|)\,,$
@@
for every $u,v\in L^{1}(\mathbb{R}^{2},\omega\,{\rm d}x)$. Concerning $A_{2}$, since $\ln(b+r)\leq r$ for every $r\geq 0$ (with the strict inequality if $r>0$), then, by Hardy-Littlewood-Sobolev inequality (Proposition 2.5), there exists a positive constant $C$ such that
@@
and we are going to prove that $\mathcal{I}$ is well-defined in $X:=D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$, see Lemma 3.2. For the rest of the paper we always use this notation to indicate our space. Before going into details of the proof, we prepose an extension of Lemma [19, Lemma 2.1] to our framework. It will be crucial in order to transfer estimates from the bilinear form $A_{1}$ to the norm $\|\cdot\|_{\ast,p}$, since it will be mainly applied with $\varphi_{n}=F(u_{n})$ in combination with the lower bound (2.8).
@@
Let $p>1$, $u\in L^{p}(\mathbb{R}^{2})\setminus\{0\}$ and nonnegative sequences $\{u_{n}\}_{n}\subset L^{p}(\mathbb{R}^{2})$ such that $u_{n}\to u$ a.e. in $\mathbb{R}^{2}$ as $n\to+\infty$, and $\{\varphi_{n}\}_{n}\subset L^{1}(\mathbb{R}^{2})$. Let moreover $F\in C(\mathbb{R})$ with $F(t)>0$ for $t>0$.
@@
Since $u_{n}\to u$ a.e. in $\mathbb{R}^{2}$ and $F$ is continuous, by Egorov’s theorem there exist $n_{0}\in\mathbb{N}$, $R>0$, and $\delta>0$, and a measurable set $A\subset B_{R}(0)$ with positive measure, such that $F(u_{n}(x))\geq\delta$ for all $n\geq n_{0}$. For $x\in A$ and $y\in\mathbb{R}^{2}\setminus B_{(1+b)R}(0)$ we have
@@
$\begin{split}A_{1}(F(u_{n}),\varphi_{n})&\geq\int_{\mathbb{R}^{2}\setminus B_{(b+1)R}(0)}\int_{A}\ln(b+|x-y|)F(u_{n}(x))\varphi_{n}(y)\,{\rm d}x\,{\rm d}y\\ &\geq C\left(\int_{A}F(u_{n}(x))\,{\rm d}x\right)\left(\int_{\mathbb{R}^{2}\setminus B_{(b+1)R}(0)}\ln\left(b(1+|y|)^{\frac{1}{b+1}}\right)\varphi_{n}(y)\,{\rm d}y\right)\\ &\geq C\frac{\delta|A|}{b+1}\int_{\mathbb{R}^{2}\setminus B_{(b+1)R}(0)}\ln\left(b^{b+1}(1+|y|)\right)\varphi_{n}(y)\,{\rm d}y\\ &\geq C\frac{\delta|A|}{b+1}\int_{\mathbb{R}^{2}\setminus B_{(b+1)R}(0)}\ln\left(b+|y|\right)\varphi_{n}(y)\,{\rm d}y\\ &\geq C\frac{\delta|A|}{b+1}\big{(}\|\varphi_{n}\|_{\ast,1}-\ln(b+(b+1)R)\|\varphi_{n}\|_{1}\big{)},\end{split}$
@@ 3.1. Regularity of the functional $\mathcal{I}$
Under ($f_{1}$) and ($f_{sc}$) or ($f_{c}$), the functionals $I_{1},I_{2},I_{0}$ and $\mathcal{I}$ are well-defined and of class $C^{1}$ on $X$, and
$\mathcal{I}^{\prime}(u)[v]=\int_{\mathbb{R}^{2}}\nabla u\cdot\nabla v+2A_{0}(F(u),f(u)v)=\int_{\mathbb{R}^{2}}\nabla u\cdot\nabla v\,{\rm d}x+\int_{\mathbb{R}^{2}}\left(\ln|\cdot|\ast F(u)\right)f(u)v\,{\rm d}x\,.$
@@
On the other hand, combining (3.5) with (2.6) or (2.7) for $r>\frac{3}{4}p>\frac{3}{2}>1$ and $\alpha>\alpha_{0}$, one has
(3.7) $\begin{split}I_{2}(u)&\leq b\|F(u)\|_{\frac{4}{3}}^{2}\lesssim\|u\|_{\frac{4}{3}p}^{2p}+\left(\int_{\mathbb{R}^{2}}|u|^{\frac{4}{3}r}{\rm e}^{\frac{4}{3}\alpha|u|^{2}}\,{\rm d}x\right)^{\frac{3}{2}}\\ &\lesssim\|u\|_{\frac{4}{3}p}^{2p}+\|\nabla u\|_{2}^{2r}\left(\int_{\mathbb{R}^{2}}\left(\frac{u}{\|\nabla u\|_{2}}\right)^{\frac{4}{3}r}{\rm e}^{\frac{4}{3}\alpha\|u\|^{2}\left(\frac{u}{\|\nabla u\|_{2}}\right)^{2}}\,{\rm d}x\right)^{\frac{3}{2}}\\ &\lesssim\|u\|^{\frac{3}{2}p}_{p}\|\nabla u\|^{\frac{p}{2}}_{2}+C(u)\|\nabla u\|_{2}^{2r-\frac{3}{2}p}\|u\|_{p}^{\frac{3}{2}p}<+\infty\,,\end{split}$
@@
$|I_{1}(u_{n})-I_{1}(u)|$ $=\left|\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln(b+|x-y|)\left(F(u_{n}(x))F(u_{n}(y))-F(u(x))F(u(y))\right)\,{\rm d}x\,{\rm d}y\right|$
@@
$\int_{\mathbb{R}^{2}}f(u_{n}(y))|v(y)|\,{\rm d}y\leq C(b)\int_{\mathbb{R}^{2}}\ln(b+|y|)f(u_{n}(y))|v(y)|\,{\rm d}y\lesssim\|f(u_{n})\|_{*,\frac{p}{p-1}}\|v\|_{\ast,p}\leq C(u)\|v\|\,,$
@@
$\begin{split}\int_{\mathbb{R}^{2}}\left|f(u_{n}(y))-f(u(y))\right||v(y)|\,{\rm d}y&\leq(\ln b)^{-1}\int_{\mathbb{R}^{2}}\ln(b+|y|)\left|f(u_{n}(y))-f(u(y))\right||v(y)|\,{\rm d}y\\ &\lesssim\|f(u_{n})-f(u)\|_{*,\frac{p}{p-1}}\|v\|_{\ast,p}=\text{o}_{n}(1)\|v\|\,,\end{split}$
@@
namely $I_{1}^{\prime}\in C(X)$. Let us now focus on $I_{2}$. For $\{u_{n}\}_{n}\subset X$ and $u\in X$ for which (3.8) hold, by (3.5) one has
@@
and tends to 00 as $n\to+\infty$, since $\|F(u_{n})-F(u)\|_{\frac{4}{3}}=\text{o}_{n}(1)$ as in (3.9), $\|F(u)\|_{\frac{4}{3}}<+\infty$ as in (3.6), and $\|F(u_{n})\|_{\frac{4}{3}}\leq C$ by continuity.
@@
$\frac{1}{2}\left|I_{2}^{\prime}(u_{n})[v]-I_{2}^{\prime}(u)[v]\right|$ $\lesssim\left(\|F(u_{n})\|_{\frac{4}{3}}\|f(u_{n})-f(u)\|_{\frac{4}{3}p^{\prime}}+\|F(u_{n})-F(u)\|_{\frac{4}{3}}\|f(u)\|_{\frac{4}{3}p^{\prime}}\right)\|v\|$
@@ 4. Analysis of Cerami sequences
which, by some compactness argument which exploits first the boundedness of such sequence in $X$, yields the existence of a weak solution. In our case, however, the proof of the boundedness of the Cerami sequence is not standard, and we need to improve the properties that such a sequence has. The abstract result contained in Proposition 2.7 allows us to find a Cerami sequence with the additional property that $\mathcal{J}(u_{n})\to 0$ as $n\to+\infty$ (see Lemma 4.2 below), where the functional $\mathcal{J}:X\to\mathbb{R}$ is given by
@@
The boundedness of $\{u_{n}\}_{n}$ in $X$ will follow then by combining (4.1) and $J(u_{n})\to 0$. This strategy, which was first employed in [43] in the high-dimensional case, was implemented in the planar case by [25, 17] in the case $f(u)=u$. Here we need to extend it to the case of a general nonlinearity. Let us first define
@@
(4.4) $I_{1}(u)=\int_{\mathbb{R}^{2}}\ln(b+|x-y|)F(u(x))F(u(y))\,{\rm d}x\,{\rm d}y\geq\ln b\left(\int_{\mathbb{R}^{2}}F(u(x))\,{\rm d}x\right)^{2}\geq C\ln b\|u\|^{2p}_{p}.$
Concerning $I_{2}$ let us refine the upper bound proved in (3.7). If $f$ satisfies assumption $(f_{sc})$ or $(f_{c})$, then by (2.6) or (2.7), respectively, one has
$\begin{split}I_{2}(u)&\lesssim\|u\|_{\frac{4}{3}p}^{2p}+\|\nabla u\|_{2}^{2r}\left(\int_{\mathbb{R}^{2}}\left(\frac{u}{\|\nabla u\|_{2}}\right)^{\frac{4}{3}r}{\rm e}^{\frac{4}{3}\alpha\|u\|^{2}\left(\frac{u}{\|\nabla u\|_{2}}\right)^{2}}\,{\rm d}x\right)^{\frac{3}{2}}\\ &\lesssim\|u\|^{\frac{3}{2}p}_{p}\|\nabla u\|^{\frac{p}{2}}_{2}+\|\nabla u\|_{2}^{2r-\frac{3}{2}p}\|u\|_{p}^{\frac{3}{2}p}\,,\end{split}$
@@
Choosing $\nu\in\left(\frac{4}{3},\frac{4}{(4-p)_{+}}\right)$, which is nonempty since $p>1$, namely
$\begin{cases}\nu>\frac{4}{3},&\text{ if }p\geq 4,\\ \nu\in\left(\frac{4}{3},\frac{4}{4-p}\right),&\text{ if }p\in(2,4),\end{cases}$
we easily infer from (4.6) that 00 is a local minimum for $\mathcal{I}$.
@@
$\begin{split}\mathcal{I}(t^{2}u(tx))&\leq\frac{t^{4}}{2}\int_{\mathbb{R}^{2}}|\nabla u(x)|^{2}\,{\rm d}x-\frac{t^{-4}}{2}\left(\ln 2+\ln t\right)\left(\int_{\mathbb{R}^{2}}F(t^{2}u(x))\,{\rm d}x\right)^{2}\\ &\leq\frac{t^{4}}{2}\int_{\mathbb{R}^{2}}|\nabla u(x)|^{2}\,{\rm d}x-\frac{\ln 2+\ln t}{2}t^{4(p-1)}\left(\int_{\mathbb{R}^{2}}|u|^{p}\,{\rm d}x\right)^{2}\to-\infty\end{split}$
@@
Let $\gamma\in\Gamma$, then $\|\nabla(\gamma(1))\|^{2}_{2}+\|\gamma(1)\|^{2}_{2}>\rho$, and by the mean value theorem there exists $\bar{t}\in[0,1]$ such that $\|\nabla(\gamma(\bar{t}))\|^{2}_{2}+\|\gamma(\bar{t})\|^{2}_{2}=\rho$. This means that $\gamma(\bar{t})=0$, hence $\mathcal{I}(\gamma(\bar{t}))\geq a_{0}$. Therefore,
@@
With the help of the abstract result Proposition 2.7, we are in a position to prove the existence of a specific Cerami sequence.
@@
$\Psi(s,v)$ $=\mathcal{I}(\rho(s,v))=\frac{1}{2}\int_{\mathbb{R}^{2}}|\nabla\rho(s,v)|^{2}\,{\rm d}x+\frac{1}{2}\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln|x-y|F(\rho(s,v)[x])F(\rho(s,v)[y])\,{\rm d}x\,{\rm d}y$
@@
that is, $\Gamma=\{\rho\circ\tilde{\gamma}:\tilde{\gamma}\in\tilde{\Gamma}\}$ and the values (4.3) and (4.9) coincide. Let now $\{\gamma_{n}\}\subset\Gamma$ be a sequence of paths such that
@@
$={\rm e}^{-6s_{n}}\int_{\mathbb{R}^{2}}|\nabla v({\rm e}^{-s_{n}}x)|^{2}\,{\rm d}x+{\rm e}^{-2s_{n}}\left(\int_{\mathbb{R}^{2}}\ln(b+|x|)|v({\rm e}^{-s_{n}}x)|^{p}\,{\rm d}x\right)^{\frac{1}{p}}$
$={\rm e}^{-4s_{n}}\int_{\mathbb{R}^{2}}|\nabla v(x^{\prime})|^{2}\,{\rm d}x^{\prime}+{\rm e}^{-2s_{n}\left(\frac{p-1}{p}\right)}\left(\int_{\mathbb{R}^{2}}\ln(b+{\rm e}^{s_{n}}|x^{\prime}|)|v(x^{\prime})|^{p}\,{\rm d}x^{\prime}\right)^{\frac{1}{p}}$
$=(1+\text{o}_{n}(1))\int_{\mathbb{R}^{2}}|\nabla v(x^{\prime})|^{2}\,{\rm d}x^{\prime}+(1+\text{o}_{n}(1))\left(\int_{\mathbb{R}^{2}}\ln(b+{\rm e}^{s_{n}}|x^{\prime}|)|v(x^{\prime})|^{p}\,{\rm d}x^{\prime}\right)^{\frac{1}{p}}$
$=(1+\text{o}_{n}(1))\|v\|^{2}+\text{o}_{n}(1)\|v\|^{p}_{p},$
@@
$\ln(b+{\rm e}^{-s_{n}}|x|)\leq\ln(b+|x|)+\ln\left(1+\frac{(1-{\rm e}^{-s_{n}})|x|}{b+|x|}\right)\leq\ln(b+|x|)+\ln(1+(1-{\rm e}^{-s_{n}}))\,.$
@@
(4.12) $\begin{split}(1+\|(s_{n},v_{n})\|_{\tilde{X}})|\Psi^{\prime}(s_{n},v_{n})(0,w_{n})|&=(1+(|s_{n}|^{2}+\|v_{n}\|^{2})^{\frac{1}{2}})|\mathcal{I}^{\prime}(\rho(s_{n},v_{n}))\rho(s_{n},w_{n})|\\ &=\left(1+\text{o}_{n}(1)+\|v_{n}\|\right)|\mathcal{I}^{\prime}(u_{n})v|\\ &=\left(1+\text{o}_{n}(1)+(1+\text{o}_{n}(1))\|u_{n}\|)\right)|\mathcal{I}^{\prime}(u_{n})v|\,,\end{split}$
@@
(4.13) $\begin{split}(1+\|(s_{n},v_{n})\|_{\tilde{X}})|\Psi^{\prime}(s_{n},v_{n})(0,w_{n})|&\leq(1+\|(s_{n},v_{n})\|_{\tilde{X}})\|\Psi^{\prime}(s_{n},v_{n})\|_{\tilde{X}^{\prime}}\|w_{n}\|\\ &=\text{o}_{n}(1)(1+\text{o}_{n}(1))\|v\|\,.\end{split}$
@@
$\left(1+\text{o}_{n}(1)+(1+\text{o}_{n}(1))\|u_{n}\|)\right)\|\mathcal{I}^{\prime}(u_{n})\|_{X^{\prime}}\to 0$
@@
for which, by (2.9), $|v_{n}|\leq(1-\tau)|u_{n}|$ hold, hence $\{v_{n}\}_{n}\subset L^{p}_{\omega}(\mathbb{R}^{2})$. Moreover, a simple computation shows that

### v2

@@ 3. The variational framework
The aim of this section is to show that $\mathcal{I}$ is indeed well-defined and regular in the space $D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$ described in Section 2. First, we state an identity which will play a crucial role throughout the paper:
@@
$(u,v)\mapsto A_{0}(u,v):=A_{1}(u,v)-A_{2}(u,v)=\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln|x-y|u(x)v(y)\,{\rm d}x\,{\rm d}y\,.$
Since $b>1$, one has
(3.3) $\ln(b+|x-y|)$ $\leq\ln(b+|x|+|y|)\leq\ln(b+b|x|+b|y|)$
$\leq\ln((b+|x|)(b+|y|)=\ln(b+|x|)+\ln(b+|y|)\,,$
@@
for every $u,v\in L^{1}(\mathbb{R}^{2},\omega\,{\rm d}x)$. Concerning $A_{2}$, since $\ln(b+r)\leq r$ for every $r\geq 0$ (with the strict inequality if $r>0$), then, by the Hardy-Littlewood-Sobolev inequality (Proposition 2.9), one has
@@
and we are going to prove that $\mathcal{I}$ is well-defined in $X:=D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$, see Proposition 3.2. For the rest of the paper we always use this notation to indicate our space. Before going into the details of the proof, we prepose an extension of [19, Lemma 2.1] to our framework. It will be crucial in order to transfer estimates from the bilinear form $A_{1}$ to the norm $\|\cdot\|_{\ast,p}$, since it will be mainly applied with $\varphi_{n}=F(u_{n})$ in combination with the lower bound (2.9).
@@
Let $p>1$, $u\in L^{p}(\mathbb{R}^{2})\setminus\{0\}$ and nonnegative sequences $\{\varphi_{n}\}_{n}\subset L^{1}(\mathbb{R}^{2})$, and $\{u_{n}\}_{n}\subset L^{p}(\mathbb{R}^{2})$ such that $u_{n}\to u$ a.e. in $\mathbb{R}^{2}$ as $n\to+\infty$. Let moreover $F\in C(\mathbb{R})$ with $F(t)>0$ for $t>0$.
@@
Since $u_{n}\to u$ a.e. in $\mathbb{R}^{2}$ and $F$ is continuous, by Egorov’s theorem there exist $n_{0}\in\mathbb{N}$, $R>0$, and $\delta>0$, and a measurable set $A\subset B_{R}$ with positive measure, such that $F(u_{n}(x))\geq\delta$ for all $n\geq n_{0}$ and $x\in A$. For $x\in A$ and $y\in\mathbb{R}^{2}\setminus B_{(1+b)R}$ we have
@@
$A_{1}(F(u_{n}),\varphi_{n})$ $\geq\int_{\mathbb{R}^{2}\setminus B_{(b+1)R}}\int_{A}\ln(b+|x-y|)F(u_{n}(x))\varphi_{n}(y)\,{\rm d}x\,{\rm d}y$
$\geq\left(\int_{A}F(u_{n}(x))\,{\rm d}x\right)\!\left(\int_{\mathbb{R}^{2}\setminus B_{(b+1)R}}\!\!\ln\left(b(1+|y|)^{\frac{1}{b+1}}\right)\varphi_{n}(y)\,{\rm d}y\right)$
$\geq\frac{\delta|A|}{b+1}\int_{\mathbb{R}^{2}\setminus B_{(b+1)R}}\ln\left(b^{b+1}(1+|y|)\right)\varphi_{n}(y)\,{\rm d}y$
$\geq\frac{\delta|A|}{b+1}\int_{\mathbb{R}^{2}\setminus B_{(b+1)R}}\ln\left(b+|y|\right)\varphi_{n}(y)\,{\rm d}y$
$\geq\frac{\delta|A|}{b+1}\big{(}\|\varphi_{n}\|_{\ast,1}-\ln(b+(b+1)R)\|\varphi_{n}\|_{1}\big{)},$
@@ 3.1. Regularity of the functional $\mathcal{I}$
Under ($f_{1}$) and ($f_{sc}$) or ($f_{c}$), the functionals $I_{1},\,I_{2},\,I_{0},$ and $\mathcal{I}$ are well-defined and of class $C^{1}$ on $X$, and moreover
$\mathcal{I}^{\prime}(u)[v]$ $=\int_{\mathbb{R}^{2}}\nabla u\cdot\nabla v\,{\rm d}x+2A_{0}(F(u),f(u)v)$
$=\int_{\mathbb{R}^{2}}\nabla u\cdot\nabla v\,{\rm d}x+\int_{\mathbb{R}^{2}}\left(\ln|\cdot|\ast F(u)\right)f(u)v\,{\rm d}x,\quad\text{ forall }u,v\in X.$
@@
On the other hand, combining (3.5) with (2.7), where $\alpha>0$ or $\alpha>\alpha_{0}$ if ($f_{sc}$) or ($f_{c}$) is assumed, respectively, and $r>\frac{3}{4}p>\frac{3}{2}>1$, one has
(3.7) $I_{2}(u)$ $\leq b\|F(u)\|_{\frac{4}{3}}^{2}\lesssim\|u\|_{\frac{4}{3}p}^{2p}+\left(\int_{\mathbb{R}^{2}}|u|^{\frac{4}{3}r}{\rm e}^{\frac{4}{3}\alpha|u|^{2}}\,{\rm d}x\right)^{\frac{3}{2}}$
$\lesssim\|u\|_{\frac{4}{3}p}^{2p}+\|\nabla u\|_{2}^{2r}\left(\int_{\mathbb{R}^{2}}\left(\frac{u}{\|\nabla u\|_{2}}\right)^{\frac{4}{3}r}{\rm e}^{\frac{4}{3}\alpha\|u\|^{2}\left(\frac{u}{\|\nabla u\|_{2}}\right)^{2}}\,{\rm d}x\right)^{\frac{3}{2}}$
$\lesssim\|u\|^{\frac{3}{2}p}_{p}\|\nabla u\|^{\frac{p}{2}}_{2}+C(u)\|\nabla u\|_{2}^{2r-\frac{3}{2}p}\|u\|_{p}^{\frac{3}{2}p}<+\infty\,,$
@@
$|I_{1}(u_{n})-I_{1}(u)|$ $=\left|\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln(b+|x-y|)\big{(}F(u_{n}(x))F(u_{n}(y))-F(u(x))F(u(y))\big{)}\,{\rm d}x\,{\rm d}y\right|$
@@
$\int_{\mathbb{R}^{2}}f(u_{n}(y))|v(y)|\,{\rm d}y$ $\leq C(b)\int_{\mathbb{R}^{2}}\ln(b+|y|)f(u_{n}(y))|v(y)|\,{\rm d}y$
$\lesssim\|f(u_{n})\|_{*,\frac{p}{p-1}}\|v\|_{\ast,p}\leq C(u)\|v\|\,,$
@@
$\int_{\mathbb{R}^{2}}\Big{|}f(u_{n}(y))-f(u(y))\Big{|}|v(y)|\,{\rm d}y&\leq(\ln b)^{-1}\int_{\mathbb{R}^{2}}\ln(b+|y|)\left|f(u_{n}(y))-f(u(y))\right||v(y)|\,{\rm d}y\\ &\lesssim\|f(u_{n})-f(u)\|_{*,\frac{p}{p-1}}\|v\|_{\ast,p}=\text{o}_{n}(1)\|v\|\,,\end{aligned}$
@@
namely $I_{1}^{\prime}\in C(X)$. Let us now focus on $I_{2}$. For $\{u_{n}\}_{n}\subset X$ and $u\in X$ for which (3.8) holds, by (3.5) one has
@@
which tends to 00 as $n\to+\infty$, since $\|F(u_{n})-F(u)\|_{\frac{4}{3}}=\text{o}_{n}(1)$ by (3.9), $\|F(u)\|_{\frac{4}{3}}<+\infty$ by (3.6), and $\|F(u_{n})\|_{\frac{4}{3}}\leq C$ by continuity.
@@
$\frac{1}{2}\left|I_{2}^{\prime}(u_{n})[v]-I_{2}^{\prime}(u)[v]\right|\lesssim\left(\|F(u_{n})\|_{\frac{4}{3}}\|f(u_{n})-f(u)\|_{\frac{4}{3}p^{\prime}}+\|F(u_{n})-F(u)\|_{\frac{4}{3}}\|f(u)\|_{\frac{4}{3}p^{\prime}}\right)\|v\|\,,$
@@ 4. Analysis of Cerami sequences
which yields the existence of a weak solution, by using some compactness argument which exploits the boundedness of such a sequence in $X$. In our case, however, the proof of the boundedness of the Cerami sequence is not standard, and we need to improve the properties that such a sequence has. The abstract result contained in Proposition 2.11 allows us to find a Cerami sequence with the additional property that $\mathcal{J}(u_{n})\to 0$ as $n\to+\infty$ (see Lemma 4.2 below), where the functional $\mathcal{J}:X\to\mathbb{R}$ is given by
@@
The boundedness of $\{u_{n}\}_{n}$ in $X$ will follow then by combining (4.1) and $\mathcal{J}(u_{n})\to 0$. This strategy, which is reminiscent of the construction of Palais-Smale-Pohožaev sequences in the context of problems with prescribed mass [30], was first employed in [44] in the higher-dimensional case, and implemented in the planar setting by [25, 17] in the case $f(u)=u$. Here we need to extend it to the case of a general nonlinearity. Let us first define
@@
(4.4) $I_{1}(u)$ $=\int_{\mathbb{R}^{2}}\ln(b+|x-y|)F(u(x))F(u(y))\,{\rm d}x\,{\rm d}y$
$\geq\ln b\left(\int_{\mathbb{R}^{2}}F(u(x))\,{\rm d}x\right)^{2}\geq C\ln b\|u\|^{2p}_{p}.$
Concerning $I_{2}$, let us refine the upper bound (3.7). By (2.7) one has
$I_{2}(u)$ $\lesssim\|u\|_{\frac{4}{3}p}^{2p}+\|\nabla u\|_{2}^{2r}\left(\int_{\mathbb{R}^{2}}\left(\frac{u}{\|\nabla u\|_{2}}\right)^{\frac{4}{3}r}{\rm e}^{\frac{4}{3}\alpha\|u\|^{2}\left(\frac{u}{\|\nabla u\|_{2}}\right)^{2}}\,{\rm d}x\right)^{\frac{3}{2}}$
$\lesssim\|u\|^{\frac{3}{2}p}_{p}\|\nabla u\|^{\frac{p}{2}}_{2}+\|\nabla u\|_{2}^{2r-\frac{3}{2}p}\|u\|_{p}^{\frac{3}{2}p}\,,$
@@
Choosing $\nu\in\left(\frac{4}{3},\frac{4}{(4-p)_{+}}\right)$, which is nonempty since $p>1$, we easily infer from (4.6) that 00 is a local minimum for $\mathcal{I}$.
@@
$\mathcal{I}(t^{2}u(tx))$ $\leq\frac{t^{4}}{2}\int_{\mathbb{R}^{2}}|\nabla u(x)|^{2}\,{\rm d}x-\frac{t^{-4}}{2}\left(\ln 2+\ln t\right)\left(\int_{\mathbb{R}^{2}}F(t^{2}u(x))\,{\rm d}x\right)^{2}$
$\leq\frac{t^{4}}{2}\int_{\mathbb{R}^{2}}|\nabla u(x)|^{2}\,{\rm d}x-\frac{\ln 2+\ln t}{2}t^{4(p-1)}\left(\int_{\mathbb{R}^{2}}|u|^{p}\,{\rm d}x\right)^{2}\to-\infty$
@@
Let $\gamma\in\Gamma$, then $\|\nabla\gamma(1)\|^{2}_{2}+\|\gamma(1)\|^{p}_{p}>\rho$, and by the mean value theorem there exists $\bar{t}\in[0,1]$ such that $\|\nabla\gamma(\bar{t})\|^{2}_{2}+\|\gamma(\bar{t})\|^{p}_{p}=\rho$. This means that $\gamma(\bar{t})\neq 0$, hence $\mathcal{I}(\gamma(\bar{t}))\geq a_{0}$. Therefore,
@@
With the help of the abstract result of Proposition 2.11, we are in a position to prove the existence of a specific Cerami sequence.
@@
$\Psi(s,v)$ $=\mathcal{I}(\rho(s,v))$
$=\frac{1}{2}\int_{\mathbb{R}^{2}}|\nabla\rho(s,v)|^{2}\,{\rm d}x+\frac{1}{2}\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln|x-y|F(\rho(s,v)[x])F(\rho(s,v)[y])\,{\rm d}x\,{\rm d}y$
@@
that is, $\Gamma=\{\rho\circ\tilde{\gamma}:\tilde{\gamma}\in\tilde{\Gamma}\}$ and the values (4.3) and (4.9) coincide. Let now $\{\gamma_{n}\}_{n}\subset\Gamma$ be a sequence of paths such that
@@
$={\rm e}^{-6s_{n}}\int_{\mathbb{R}^{2}}|\nabla v({\rm e}^{-s_{n}}x)|^{2}\,{\rm d}x+{\rm e}^{-2s_{n}}\left(\int_{\mathbb{R}^{2}}\ln(b+|x|)|v({\rm e}^{-s_{n}}x)|^{p}\,{\rm d}x\right)^{\frac{2}{p}}$
$={\rm e}^{-4s_{n}}\int_{\mathbb{R}^{2}}|\nabla v(x^{\prime})|^{2}\,{\rm d}x^{\prime}+{\rm e}^{-2s_{n}\left(\frac{p-1}{p}\right)}\left(\int_{\mathbb{R}^{2}}\!\ln(b+{\rm e}^{s_{n}}|x^{\prime}|)|v(x^{\prime})|^{p}\,{\rm d}x^{\prime}\right)^{\frac{2}{p}}$
$=(1+\text{o}_{n}(1))\left(\int_{\mathbb{R}^{2}}|\nabla v(x^{\prime})|^{2}\,{\rm d}x^{\prime}+\left(\int_{\mathbb{R}^{2}}\!\ln(b+{\rm e}^{s_{n}}|x^{\prime}|)|v(x^{\prime})|^{p}\,{\rm d}x^{\prime}\right)^{\frac{2}{p}}\right)$
$=(1+\text{o}_{n}(1))\|v\|^{2}+\text{o}_{n}(1)\|v\|^{2}_{p},$
@@
$\ln(b+{\rm e}^{-s_{n}}|x|)$ $\leq\ln(b+|x|)+\ln\left(1+\frac{(1-{\rm e}^{-s_{n}})|x|}{b+|x|}\right)$
$\leq\ln(b+|x|)+\ln(1+(1-{\rm e}^{-s_{n}}))\,.$
@@
(4.12) $(1+\|(s_{n},v_{n})\|_{\tilde{X}})|\Psi^{\prime}(s_{n},v_{n})(0,w_{n})|$ $=(1+(|s_{n}|^{2}+\|v_{n}\|^{2})^{\frac{1}{2}})|\mathcal{I}^{\prime}(\rho(s_{n},v_{n}))\rho(s_{n},w_{n})|$
$=\left(1+\text{o}_{n}(1)+(1+\text{o}_{n}(1))\|u_{n}\|)\right)|\mathcal{I}^{\prime}(u_{n})v|\,,$
@@
(4.13) $(1+\|(s_{n},v_{n})\|_{\tilde{X}})|\Psi^{\prime}(s_{n},v_{n})(0,w_{n})|$ $\leq(1+\|(s_{n},v_{n})\|_{\tilde{X}})\|\Psi^{\prime}(s_{n},v_{n})\|_{\tilde{X}^{\prime}}\|w_{n}\|$
$=\text{o}_{n}(1)(1+\text{o}_{n}(1))\|v\|\,.$
@@
$\left(1+\text{o}_{n}(1)+(1+\text{o}_{n}(1))\|u_{n}\|\right)\|\mathcal{I}^{\prime}(u_{n})\|_{X^{\prime}}\to 0$
@@
where $\tau$ is defined in $(f_{2})$, and for which, by (2.10), $|v_{n}|\leq(1-\tau)|u_{n}|$ hold, hence $\{v_{n}\}_{n}\subset L^{p}_{\omega}(\mathbb{R}^{2})$. Moreover, a simple computation shows that